wiki:UC20191007

Version 1 (modified by trac, 16 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 10/07/19

Attending: Kevin Schlaufman (JHU), Ben Williams (UW), Mukremin Kilic (OU), Anne Verbiscer (UVa), John Bally (CU), Sarah Tuttle (UW), Nancy Chanover (NMSU), Derek Buzasi (FGCU), Bill Ketzeback (APO), Moire Prescott (NMSU), Eric Bellm (UW), Misty Bentz (GSU), John Wilson (UVa)


User feedback and comments from institutional representatives:

  • Wyoming - no report
  • Washington - nothing to report
  • UVa - two items from Steve Majewski:1) he was invited to participate in discussions with WIYN about potential time trades among APO/WIYN/SOAR, and (2) he recommends keeping ARCSAT on the SkyNet network. ARCSAT will be discussed in more detail below. Nancy and Steve are coordinating re: ARC position's on time trades (we are in favor of it!) and will continue to work this with WIYN.
  • Seattle - no report
  • Oklahoma - nothing to report
  • NMSU - nothing to report
  • NAPG - no report
  • JHU - nothing to report
  • Georgia State - nothing to report
  • FGCU - nothing to report
  • Colorado - nothing to report

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 8/01/2019 - 8/29/2019

0) Overview

We had minimal problems this period. It was a wet month but not a total shut out.

1) Telescope

Nominal operations with minimal problems.

2) Instruments

DIS: The DIS blue camera continues to see contamination issues It was vacuum serviced but as expected deteriorated quickly . Red camera has stabilized and may start seeing improvement with cooling temperatures, however the ion pump is showing signs of aging and is shutting down frequently. We could service it this month but only if really necessary.

TripleSpec is operational.

Agile is operational.

Echelle is operational. The IOL is steadily getting worse, we are monitoring it closely to determine when and IF we will intervene. The blue value has gotten worse than our nominal limit and red is stable. We expect blue will recover some with cooler temperatures so we are watching it regularly. Let us know if it will affect your science.

NIC-FPS is operational.

ARCTIC is operational.

3) ARCSAT

Operating nominally. The new guide scope mount is installed, aligned and will be operational shortly after engineering this week. The visible light cloud camera got a makeover.


Additional telescope and instrument discussion:

September was a pretty wet month (it was often humid if not raining), which resulted in a fair amount of observing time lost due to weather. The telescope has been performing nominally; there were recently some issues related to the altitude axis (likely due to recent temperature changes); the engineering staff will look at it more seriously over the next week, particularly if the problem persists or gets worse. The obs-specs should be able to spot this right away if it recurs, but if observers notice that the telescope stops tracking and see their objects trailing off the guider or imaging field of view, please report it immediately.

TripleSpec, Agile, ARCTIC and NIC-FPS have all been performing well with no or very minor issues. The echelle is also performing well at the moment; the long-standing inter-order light issue has improved somewhat as the temperatures have cooled. Bill expects the performance at the blue end to continue to improve until about March and then it may turn back around and get worse again. To recap the situation, this is likely a vacuum issue where we are getting water vapor condensing in front of the CCD. It is a longstanding problem that we have been monitoring for more than 7 years. It is similar to what we see in DIS, but since ARCES is such a big assembly, trying to track down the cause and replacing O-rings would be a monumental task that would take the instrument off-line for an extended period of time. The DIS blue contamination is very bad - we vacuum serviced it over the past month and subsequently the performance was quite good for 3-4 hrs afterwards and then it quickly deteriorated. DIS red has degraded a little bit compared to where it had been over the last few months; its ion pump is showing its age and likely will need to be replaced over the next few months.

ARCSAT has been performing fine. We installed the new mount for the off-axis guide scope, and it was aligned with the science cameras. We still need to do some engineering adjustments to the software configurations for guiding; we hope to get it done this week. A lightning strike last week may have taken out a few things in the ARCSAT dome; this will be investigated this week. The telescope, cameras, and network all seem OK, but we can’t talk to some of the weather metrology and we have lost communications with the dome, which made the telescope unavailable over the past weekend.

Some additional discussion of ARCSAT ensued. Kevin said that he talked to Adam Riess about the potential for low-cost IR cameras on small telescopes. There have been some recent developments in those directions using cheap detectors that Rob Simcoe at MIT has demonstrated, using them for bright things. These are InGaAs IR detectors. Could be a low cost way to build an IR imaging (J & H band only) instrument for a small telescope that could be used for looking at bright things. These detectors are cheaper by about an order of magnitude than traditional FPAs; if we need a wide FOV perhaps they could be mosaicked together. The paper from Rob Simcoe’s group was recently published in AJ if people want to learn more. [Editor's note: here is the link: http://10.3847/1538-3881/aae094.] Misty shared the GSU perspective, which is that the 3.5m telescope and instrumentation is their priority, so although ARCSAT is nice to have, if funding is limited they would not want to divert resources from the 3.5m for new instrumentation for ARCSAT. Nancy thanked everyone for their input on this topic and reported that ARCSAT will a topic of discussion at the Board of Governor's meeting next month. She will present the BoG with a menu of options and their costs (including those discussed at the strategic planning meeting in June) in an effort to inform the Board's discussion regarding whether ARCSAT needs its own budget line and how it should be handled going forward.


Q4 3.5m and ARCSAT scheduling:

Both the 3.5m and ARCSAT schedules for Q4 are out. The Q4 3.5m schedule is pretty packed; there is not much OPEN time and we’ve had a lot of requests for the time we do have. Russet managed to arrange quite a few class trips that were challenging to schedule - she worked VERY hard to assemble the complicated Q4 schedule, and she should be commended and thanked! The Q4 ARCSAT schedule is also pretty full, although there are several open weeks (2 late in the quarter and also next week). Anyone wishing to request this open time should send proposals to Russet, Ben, and Nancy. Russet also worked on two additional changes to the published version of the 3.5m observing schedule. The format for the online 3.5m schedule has been modified so we can now readily see if a given program is a ToO or not, or if it is protected from unscheduled interruptions. The format of the night logs has also changed; it now includes ToO status and is easier to parse and extract statistics about telescope usage.


NA2 Baffle Project:

This project is moving forward towards completion. We held a pre-fabrication review last Thursday afternoon, and we now have a fully designed baffle system for NA2 that was analyzed for stray light by Photon Engineering. The new baffle will yield an 8’ x 8’ FOV (the old baffle provided a 6.1’ diameter (round) FOV). The metric that was studied in detail was the degree of stray light relative to original baffle; when one opens up the FOV like this, the focal plane sees more parts of the telescope that can be illuminated by more sources on the sky. Unsurprisingly, with the more open baffle we do see a small increase in stray light that comes from far off-axis. This means that if one is observing a target with ARCTIC and there is a bright Moon 20 degrees away, there will be a slightly higher background level in the images. This increase did not vary much between the multiple baffle designs that we analyzed. If this is an issue for people we need to know - we can send plots that describe this in more detail. The second major item of discussion was the guider illumination pattern. Increasing the footprint of the baffle means that there is less light going to guider, but per Bill's design the square baffle box has a slot in it, which increases light to guider and makes the guider more uniformly illuminated relative to current performance. They have a fabrication schedule with shop drawings that are ready to go. It can likely be fabricated by late December, with commissioning taking place in late January. Ben asked how difficult it would be to switch between baffles and go back to the previous one? Answer: it can be done during the day in a matter of 15-30 minutes; it remains to be seen how easy it would be to do in the middle of the night since it will likely be a two-person job. The new baffle weighs roughly 1.7x the weight of the old one. Muk asked if we could quantify how much additional scattered light we would get, and whether we can provide users with that information ahead of time? Answer: we already have the point source transmittance plots that describe how much sensitivity you have as you go towards 45 degrees off-axis, and we can certainly share those. When the new NA2 baffle is commissioned on the telescope we will be taking before-and-after data so that users can assess the difference using actual observations.


KOSMOS Update:

Preparations for delivering KOSMOS II to APO are continuing. The UW TEG received the old cable (that had been rewired) from APO; next they will be doing some signal tests with everything installed into old dewar, and then moving it to new dewar. They were pumping down the dewar over the weekend and will be ready to cool it tomorrow. The slit viewer is ready and waiting. Kal is finalizing the electronics in second electronics box, which controls the internal calibration lamps and slit viewer. So far they are on schedule. Conor will try out the software when they bring up the old dewar. CTIO is still using COSMOS so by making the changes we did on KOSMOS to accommodate the larger (COSMOS) dewar, that has allowed us to have a true sibling instrument with the other version that is still being actively used at CTIO. Sarah expressed some concern about how the changes in background light with the new NA2 baffle will affect KOSMOS, but we'll see. If people have desired sources that they want us to look at during the KOSMOS II commissioning at APO, notify your Users Committee reps; we will discuss this again on the November UC telecon. Moire asked about multi-slit capability; Sarah replied that those interested in multi-slit options should send her an email so they can discuss how to implement that. John Bally asked about the possibility of getting additional gratings. Nancy replied that this is still on her radar but we likely won't order anything until we have demonstrated what we have and what the performance looks like during and after commissioning.


Status Update on Near-Term Action Items from Strategic Planning Meeting

  • DAWG: the DAWG is meeting approximately monthly. Their first initiative is to assess the tools, pipelines, and software packages that people are using to reduce their APO data. They are collecting this info through a survey; all APO users are encouraged to complete the survey: https://docs.google.com/forms/d/e/1FAIpQLSdvBV1Z2O-mU9-XafRFJV4jRpWAA2Rhjhu7dT6wBTli3JHJ7g/viewform Deadline is Friday November 1!
  • Getting narrow-band filters for ARCTIC: Nancy sent out a survey to 35m-general soliciting input concerning desired narrow-band filters that can take advantage of the full (soon-to-be-unvignetted) FOV of ARCTIC, as well as additional desired gratings for KOSMOS. The results are pretty mixed (there was not an overwhelming outcry for one or two specific filters), so Nancy will discuss this with the management team and make a decision so she can get them ordered before the end of CY19.
  • Science communications intern: Nancy is still awaiting approval for this internship position from the NMSU internship office. [Note added after call: it has been approved! Once NMSU's HR office approves it we can advertise; hopefully before the end of this week.]
  • AAS presence: Nancy reserved an exhibitor booth, and needs to know who from ARC and leasing institutions is a) going to the AAS meeting in January and b) willing to help out by spending some time at the booth talking about APO to passers-by. Nancy asked if there was interest in us submitting an abstract for a poster for the meeting, with the goal of increasing the visibility of APO, promoting our science, and recruiting new ARC and/or leasing partners. Sarah was agnostic about it; Derek thought it was a good idea. Nancy will see if she can get something together in time for the submission deadline of October 8.
  • Student training videos: There were several good ideas that were brainstormed during last month's call. Nancy will contact trip leaders to see what might be possible.

Request for Science Highlights for BoG report

Nancy is requesting science highlights from each ARC member institution and leasing partner to present at the ARC Board of Governors meeting in early November. The deadline for submitting them is October 15. UC reps are asked to collect these from their users and send them to Nancy or have them sent directly. There is no template to follow but the slides should include project title, PI and team member names, institution, a graphic, some bullet points summarizing the take-away result, and journal references (if available).


ACTION ITEMS:

Open action items from previous meetings:

  • Ben will send out a link to KOSMOS info for interested prospective users. This is intended to be a page where prospective users could learn about the general capabilities of KOSMOS at a glance. STATUS: CLOSED. This information is now posted on the users' wiki.
  • Nancy will coordinate with Mike Crenshaw regarding distribution of the Astro2020 APC White Paper. STATUS: CLOSED. It is posted on the users' wiki, and Nancy will alert the Board to its existence when she gives her annual report.
  • UC reps should remind their users to complete the survey about desired narrow-band ARCTIC filters and KOSMOS gratings by the deadline of September 20. STATUS: CLOSED. The survey deadline has passed; 10 people submitted input.
  • Nancy will contact trip leaders bringing student groups to APO in Q4 to discuss training video ideas. STATUS: OPEN.
  • UC reps should poll their users to see who is going to the AAS meeting and would be willing to help out with booth presence. STATUS: OPEN.
  • UC reps should request science highlights from their users and have them sent to Nancy no later than October 15. STATUS: OPEN.

New action items from this meeting:

None.


All Other Business

None.


Next meeting:

The next Users Committee telecon will be on Tuesday November 5 at 10:30 am MST.


Note: See TracWiki for help on using the wiki.